A 400 mL buffer solution was prepared by reacting sodium acetate, CH3COONa,

with 1.00 M acetic acid, CH3COOH, with a pH of 4.56 Calculate the mass of

CH3COONa required to prepare the solution. Ka CH3COOH= 1.75 x 10-5


1
Expert's answer
2022-06-09T21:29:03-0400

1.

Using pH, we find the concentration of H + ions

pH = 4,56

pH= -log [H+]

4,56 = -log [H+]

[H+] = 2,76 * 10-5

2.

Using the H + ion concentration and the Ka value, we find the CH3COOH concentration. for this we divide the H + ion concentration by Ka:

CH3COOH => CH3COO– + H+

[CH3COOH] = [H+] /Ka =2,76*10-5/1,75*10-5 =

= 1,577 mol/l

3.

CH3COONa + H2O = CH3COOH + NaOH

We find the amount of CH3COONa for this we separate the CH3COOH concentration given in the task from the CH3COOH concentration found Then multiply it by 0.4 l

1,577 M - 1 M = 0,577 M

n= CM * V = 0,577 * 0,4 = 0,2308 mol

4.

Mr (CH3COONa) = 82 g/mol

We find the mass of CH3COONa using the following formula:

m = n * Mr = 0,2308 * 82 = 18,9 g


Answer: 18,9 g CH3COONa